# Explain Yourself, Briefly! Self-Explaining Neural Networks with Concise Sufficient Reasons

### v2

@@ 1 Introduction
Figure 1: An example of a sufficient reason generated by a model trained with sufficient subset training (SST) on the IMAGENET dataset, compared to those generated by post-hoc methods on standard-trained models. While explanations from Anchors and GS are larger, those from SIS are less faithful and lack subset sufficiency (details in the experiments in Section 5). SST generates explanations that are both concise and faithful, while performing this task with significantly improved efficiency. Additional examples appear in appendix F.
@@ 3 Problems with post-hoc minimal sufficient reason computations
The highly intractable complexity class $\Sigma^{P}_{2}$ encompasses problems solvable in NP time with constant-time access to a coNP oracle, whereas $\text{NP}^{\text{PP}}$ includes problems solvable in NP time with constant-time access to a probabilistic PP oracle. More details on these classes are available in Appendix A. For our purposes, it suffices to understand that NP $\subsetneq\Sigma^{P}_{2}$ and NP $\subsetneq\text{NP}^{\text{PP}}$ is widely accepted (Arora & Barak (2009)), highlighting their substantial intractability (problems that are significantly harder than NP-Complete problems).
Proof sketch. The full proof is relegated to Appendix B. We begin by extending the complexity results for robust and probabilistic sufficient reasons provided by (Barceló et al. (2020); Wäldchen et al. (2021)), which focused on either CNF classifiers or multi-layer perceptrons with binary inputs and outputs. We provide complexity proofs for a more general setting considering neural networks over any discrete or continuous input or output domains. Membership outcomes to continuous domains stem from realizing that instead of guessing certificate assignments for binary inputs, one can guess the activation status of any ReLU constraint. For hardness, we perform a technical reduction that constructs a neural network with a strictly stronger subset sufficiency. We also provide a novel complexity proof, which indicates the intractability of the relaxed version of computing cardinally minimal baseline sufficient reasons, via a reduction from the classic CNF-SAT problem.
To further emphasize the hardness of the previous results, we prove that their approximation remains intractable. We establish this through approximation-preserving reductions from the Shortest-Implicant-Core (Umans (1999)) and Max-Clique (Håstad (1999)) problems (proofs provided in Appendix C):
@@
Given a neural network classifier $f$ with ReLU activations and $\textbf{x}\in\mathbb{R}^{n}$, (i) $\forall\epsilon>0$approximating cardinally minimal robust sufficient reasons with $n^{\frac{1}{2}-\epsilon}$ factor is $\Sigma^{P}_{2}$-Hard, (ii) $\forall\epsilon>0$approximating cardinally minimal probabilistic or baseline sufficient reasons with factor $n^{1-\epsilon}$ is NP-Hard.
@@ 4.2 Masking
    Probabilistic Masking. Given some distribution $\mathcal{D}$, we sample the corresponding assignments from $\mathcal{D}(\textbf{z}|\textbf{z}_{S}=\textbf{x}_{S})$ and propagate $h_{1}(\textbf{x}_{S};\textbf{z}_{\bar{S}})$, optimizing $S$ to constitute as a probablistic sufficient reason concerning $\mathcal{D}$.
@@ 5.1 Sufficient Subset Training for Images
We train SST-based models on three prominent image classification tasks: MNIST (Deng (2012)) digit recognition using a feed-forward neural network, CIFAR-10 (Krizhevsky et al. (2009)) using Resnet18 (He et al. (2016)), and IMAGENET using a pre-trained Resnet50 model (He et al. (2016)). To simplify the training process, we set the threshold at $\tau=0.5$ and the faithfulness coefficient at $\lambda=1$, conducting the grid search exclusively over the cardinality coefficient $\xi$. Full training details can be found in Appendix D. We assess the inherent sufficient reasons generated by SST and compare them with those from standard-trained models using post-hoc methods: Anchors (Ribeiro et al. (2018)), SIS (Carter et al. (2019)), and a vision-based technique (Fong & Vedaldi (2017)) which we term gradient search (GS).
Robust Sufficient Reasons in Images. In image classification tasks, it is common to evaluate the robust sufficiency of the identified sufficient reasons, by ensuring they remain sufficient across a continuous domain (Wu et al. (2024b); Ignatiev et al. (2019); Bassan & Katz (2023)). We aim to explore the ability of SST-based models to generate such subsets, initially training all models using robust masking with a PGD $\ell_{\infty}$ ($\epsilon=0.12$) attack on the features in $\bar{S}$, as described in section 4.2. Further details on the selection of $\epsilon=0.12$ can be found in Appendix D. We subsequently assess these results against subsets produced by the aforementioned post-hoc methods over standard-trained models (i.e., with only a prediction output, and no explanation) and examine their capacity to uphold robust sufficiency. Figure 3 illustrates visualizations of these results, while Table 1 provides a summary of the empirical evaluations.
@@
Figure 3: Examples of sufficient reasons produced by SST compared to the ones generated by post-hoc approaches for MNIST, CIFAR-10, and IMAGENET. Additional examples appear in appendix F.
@@
Faithfulness-Cardinality Trade-off. Optimizing models trained with SST involves a trade-off between two primary factors: (i) the size of the generated subset and (ii) the faithfulness of the model. For instance, a full set explanation ($S:=\{1,\ldots,n\}$) scores low in terms of the optimization for small subset sizes but guarantees perfect faithfulness. Conversely, an empty set explanation ($S:=\emptyset$) achieves the best score for subset size due to its minimal cardinality but generally exhibits poor faithfulness. In our experiment, we aim to illustrate this trade-off, which can be balanced by altering the cardinality loss coefficient $\xi$. Greater $\xi$ values emphasize the subset size loss component, leading to smaller subsets at the cost of their faithfulness. Conversely, smaller $\xi$ values reduce the component’s impact, resulting in larger subsets with better faithfulness. Figure 4 provides a comparative analysis of this dynamic using the MNIST dataset with baseline masking. For greater $\xi$ values, the cardinality of the mask is maximal. As $\xi\to 0$, the explanation size converges to $50\%$ of the input. We believe this happens because the threshold is set to the default value of $\tau:=0.5$ (see Appendix D for more details), resulting in the arbitrary selection of a $50\%$ sufficient subset due to random initialization at the beginning of training. This subset is likely sufficient to determine the prediction on its own (as demonstrated in Table 2, even $23.69\%$ of the input may be sufficient for a prediction). Consequently, as the cardinality loss coefficient gets very close to zero, the parameters of the explanation component (i.e., $h_{2}$, as discussed in Section 4) remain nearly unchanged throughout training, with the explanation size fixed at $50\%$.
@@ 5.2 Sufficient Subset Training for Language
Figure 5: Explanations generated by SST using baseline vs. probabilistic masking. When each token in the complement $\overline{S}$ is replaced with the MASK token, the prediction stays negative. In the probabilistic setting, the prediction remains negative when values from $\overline{S}$ are randomly sampled. Further examples can be found in Appendix F.
@@ 6 Related Work
Post-hoc sufficient reason techniques. Some common post-hoc approaches include Anchors (Ribeiro et al. (2018)), or SIS (Carter et al. (2019; 2021)), and typically seek to obtain concise subsets which uphold probablistic sufficiency (Wang et al. (2021a); Ribeiro et al. (2018); Blanc et al. (2021)) or baseline sufficiency (Hase et al. (2021); Chockler et al. (2021); DeYoung et al. (2019); Chockler & Halpern (2024)). A different set of works is dedicated to obtaining robust sufficient reasons, where sufficiency is determined for some entire discrete or continuous domain. These are often termed abductive explanations (Ignatiev et al. (2019); Bassan et al. (2023); Ordyniak et al. (2023)) or prime implicants (Ignatiev et al. (2019); Shih et al. (2018)). Due to the complexity of producing such explanations, they are commonly obtained on simpler model types, such as decision trees (Izza et al. (2020); Huang et al. (2021); Bounia & Koriche (2023)), KNNs (Barceló et al., 2025), tree ensembles (Izza & Marques-Silva (2021); Ignatiev et al. (2022); Audemard et al. (2022b; a); Boumazouza et al. (2021)), linear models (Marques-Silva et al. (2020)), or small-scale neural networks (Ignatiev et al. (2019); Wu et al. (2024b); La Malfa et al. (2021); Bassan & Katz (2023)), typically obtained with the use of neural network verifiers (Wang et al., 2021b; Wu et al., 2024a; Fel et al., 2023).
@@
Self-explaining neural networks. Unlike post-hoc methods, some techniques modify training to improve explanations (Ismail et al. (2021); Chen et al. (2019b); Hase et al. (2021); Vafa et al. (2021); Yan & Wang (2023)) or develop a self explaining neural network (SENN) (Alvarez Melis & Jaakkola, 2018) architecture that inherently provides interpretations for its decisions (Lee et al., 2022; Shwartz et al., 2020; Rajagopal et al., 2021; Guyomard et al., 2022; Guo et al., 2023; Zhang et al., 2022). Self-explaining models typically provide inherent additive feature attributions, assigning importance weights to individual or high-level features. For example,  Chen et al. (2019a) and  Wang & Wang (2021) describe model outputs by comparing them to relevant “prototypes”, while Alvarez Melis & Jaakkola (2018) derive concept coefficients from feature transformations. Other approaches, like those by Agarwal et al. (2021) and Jain et al. (2020), focus on feature-specific neural networks or apply classifiers to snippets for NLP explanations, respectively.  Koh et al. (2020) predict outcomes based on high-level concepts.
Due to their focus on additive attributions, these methods generally operate under the implicit or explicit assumption that the model’s behavior can be approximated as linear under certain conditions (Yeh et al. (2019); Lundberg & Lee (2017)). This assumption is crucial as it underpins the approach of decomposing a model’s output into independent, additive contributions from individual features, and may hence overlook nonlinear interactions among features (Slack et al. (2020); Ribeiro et al. (2018); Yeh et al. (2019)). In contrast, we propose a self-explaining framework that provides a distinct form of explanation — sufficient reasons for predictions — which do not depend on assumptions of underlying linearity (Ribeiro et al. (2018)). To our knowledge, this is the first self-explaining framework designed to generate such explanations.
@@ 7 Limitations
A primary limitation of SST is the need to tune additional hyperparameters due to multiple optimization objectives. We simplified this by conducting a grid search solely over the cardinality loss coefficient $\xi$, as detailed in Appendix D. Additionally, the dual-propagation training phase is twice more computationally demanding than standard training, and this complexity escalates with challenging masking strategies, such as robust masking (see Appendix E for more details). However, this investment in training reduces the need for complex post-training computations, as highlighted in Section 5. Lastly, overly optimizing our interpretability constraints can sometimes reduce predictive accuracy. Our experiments showed this was true for IMAGENET, though all other benchmarks retained comparable accuracy levels (within $2\%$, and under $1\%$ for all language-model benchmarks). Interestingly, accuracy even improved under some configurations for MNIST classification. Future research endeavors could further investigate the conditions under which SST can be utilized not only for its inherent explanations but also to enhance predictive accuracy.
@@ Acknowledgments
  • Appendix A contains background for the computational complexity proofs.
  • Appendix B contains the proof of theorem 1.
  • Appendix C contains the proof of theorem ii.
  • Appendix D contains technical specifications, related to the models, and training.
  • Appendix E contains information regarding the training time of SST compared to standard training.
  • Appendix F contains supplementary results.
Appendix A Computational Complexity Background

### v3

@@ 1 Introduction
Figure 1: An example of a sufficient reason generated by a model trained with sufficient subset training (SST) on the IMAGENET dataset, compared to those generated by post-hoc methods on standard-trained models. While explanations from Anchors and GS are larger, those from SIS are less faithful and lack subset sufficiency (details in the experiments in Section 5). SST generates explanations that are both concise and faithful, while performing this task with significantly improved efficiency. Additional examples appear in appendix J.
@@ 3 Problems with post-hoc minimal sufficient reason computations
The highly intractable complexity class $\Sigma^{P}_{2}$ encompasses problems solvable in NP time with constant-time access to a coNP oracle, whereas $\text{NP}^{\text{PP}}$ includes problems solvable in NP time with constant-time access to a probabilistic PP oracle. More details on these classes are available in Appendix B. For our purposes, it suffices to understand that NP $\subsetneq\Sigma^{P}_{2}$ and NP $\subsetneq\text{NP}^{\text{PP}}$ is widely accepted (Arora & Barak (2009)), highlighting their substantial intractability (problems that are significantly harder than NP-Complete problems).
Proof sketch. The full proof is relegated to Appendix C. We begin by extending the complexity results for robust and probabilistic sufficient reasons provided by (Barceló et al. (2020); Wäldchen et al. (2021)), which focused on either CNF classifiers or multi-layer perceptrons with binary inputs and outputs. We provide complexity proofs for a more general setting considering neural networks over any discrete or continuous input or output domains. Membership outcomes to continuous domains stem from realizing that instead of guessing certificate assignments for binary inputs, one can guess the activation status of any ReLU constraint. For hardness, we perform a technical reduction that constructs a neural network with a strictly stronger subset sufficiency. We also provide a novel complexity proof, which indicates the intractability of the relaxed version of computing cardinally minimal baseline sufficient reasons, via a reduction from the classic CNF-SAT problem.
To further emphasize the hardness of the previous results, we prove that their approximation remains intractable. We establish this through approximation-preserving reductions from the Shortest-Implicant-Core (Umans (1999)) and Max-Clique (Håstad (1999)) problems (proofs provided in Appendix D):
@@
Given a neural network $f$ with ReLU activations and $\textbf{x}\in\mathbb{R}^{n}$, (i) $\forall\epsilon>0$approximating cardinally minimal robust sufficient reasons with $n^{\frac{1}{2}-\epsilon}$ factor is $\Sigma^{P}_{2}$-Hard, (ii) $\forall\epsilon>0$approximating cardinally minimal probabilistic or baseline sufficient reasons with factor $n^{1-\epsilon}$ is NP-Hard.
@@ 4.2 Masking
    Probabilistic Masking. Given some distribution $\mathcal{D}$, we sample the corresponding assignments from $\mathcal{D}(\textbf{z}|\textbf{z}_{S}=\textbf{x}_{S})$ and propagate $h_{1}(\textbf{x}_{S};\textbf{z}_{\bar{S}})$, optimizing $S$ to constitute as a probabilistic sufficient reason concerning $\mathcal{D}$.
@@ 5.1 Sufficient Subset Training for Images
We train SST-based models on three prominent image classification tasks: MNIST (Deng (2012)) digit recognition using a feed-forward neural network, CIFAR-10 (Krizhevsky et al. (2009)) using Resnet18 (He et al. (2016)), and IMAGENET using a pre-trained Resnet50 model (He et al. (2016)). To simplify the training process, we set the threshold at $\tau=0.5$ and the faithfulness coefficient at $\lambda=1$, conducting the grid search exclusively over the cardinality coefficient $\xi$. Full training details can be found in Appendix E. We assess the inherent sufficient reasons generated by SST and compare them with those from standard-trained models using post-hoc methods: Anchors (Ribeiro et al. (2018)), SIS (Carter et al. (2019)), and a vision-based technique (Fong & Vedaldi (2017)) which we term gradient search (GS).
Robust Sufficient Reasons in Images. In image classification tasks, it is common to evaluate the robust sufficiency of the identified sufficient reasons, by ensuring they remain sufficient across a continuous domain (Wu et al. (2024b); Ignatiev et al. (2019); Bassan & Katz (2023)). We aim to explore the ability of SST-based models to generate such subsets, initially training all models using robust masking with a PGD $\ell_{\infty}$ ($\epsilon=0.12$) attack on the features in $\bar{S}$, as described in section 4.2. Further details on the selection of $\epsilon=0.12$ can be found in Appendix E. We subsequently assess these results against subsets produced by the aforementioned post-hoc methods over standard-trained models (i.e., with only a prediction output, and no explanation) and examine their capacity to uphold robust sufficiency. Figure 3 illustrates visualizations of these results, while Table 1 provides a summary of the empirical evaluations.
@@
Figure 3: Examples of sufficient reasons produced by SST compared to the ones generated by post-hoc approaches for MNIST, CIFAR-10, and IMAGENET. Additional examples appear in appendix J.
@@
Faithfulness-Cardinality Trade-off. Optimizing models trained with SST involves a trade-off between two primary factors: (i) the size of the generated subset and (ii) the faithfulness of the model. For instance, a full set explanation ($S:=\{1,\ldots,n\}$) scores low in terms of the optimization for small subset sizes but guarantees perfect faithfulness. Conversely, an empty set explanation ($S:=\emptyset$) achieves the best score for subset size due to its minimal cardinality but generally exhibits poor faithfulness. In our experiment, we aim to illustrate this trade-off, which can be balanced by altering the cardinality loss coefficient $\xi$. Greater $\xi$ values emphasize the subset size loss component, leading to smaller subsets at the cost of their faithfulness. Conversely, smaller $\xi$ values reduce the component’s impact, resulting in larger subsets with better faithfulness. Figure 4 provides a comparative analysis of this dynamic using the MNIST dataset with baseline masking. For greater $\xi$ values, the cardinality of the mask is maximal. As $\xi\to 0$, the explanation size converges to $50\%$ of the input. We believe this happens because the threshold is set to the default value of $\tau:=0.5$ (see Appendix E for more details), resulting in the arbitrary selection of a $50\%$ sufficient subset due to random initialization at the beginning of training. This subset is likely sufficient to determine the prediction on its own (as demonstrated in Table 2, even $23.69\%$ of the input may be sufficient for a prediction). Consequently, as the cardinality loss coefficient gets very close to zero, the parameters of the explanation component (i.e., $h_{2}$, as discussed in Section 4) remain nearly unchanged throughout training, with the explanation size fixed at $50\%$.
@@ 5.2 Sufficient Subset Training for Language
Figure 5: Explanations generated by SST using baseline vs. probabilistic masking. When each token in the complement $\overline{S}$ is replaced with the MASK token, the prediction stays negative. In the probabilistic setting, the prediction remains negative when values from $\overline{S}$ are randomly sampled. Further examples can be found in Appendix J.
@@ 6 Related Work
Post-hoc sufficient reason techniques. Some common post-hoc approaches include Anchors (Ribeiro et al. (2018)), or SIS (Carter et al. (2019)), and typically seek to obtain concise subsets which uphold probabilistic sufficiency (Wang et al. (2021a); Ribeiro et al. (2018); Blanc et al. (2021)) or baseline sufficiency (Hase et al. (2021); Chockler et al. (2021); DeYoung et al. (2019)). A different set of works is dedicated to obtaining robust sufficient reasons, where sufficiency is determined for some entire discrete or continuous domain. These are often termed abductive explanations (Ignatiev et al. (2019); Bassan et al. (2023)) or prime implicants (Ignatiev et al. (2019); Shih et al. (2018)). Due to the complexity of producing such explanations, they are commonly obtained on simpler model types, such as decision trees (Izza et al. (2020)), tree ensembles (Izza & Marques-Silva (2021); Ignatiev et al. (2022)), linear models (Marques-Silva et al. (2020)), or small-scale neural networks (Ignatiev et al. (2019); La Malfa et al. (2021); Bassan & Katz (2023)).
@@
Self-explaining neural networks. Unlike post-hoc methods, some techniques modify training to improve explanations (Ismail et al. (2021); Chen et al. (2019b); Hase et al. (2021); Vafa et al. (2021); Yan & Wang (2023)) or develop self-explaining models that inherently provide interpretations for their decisions. Self-explaining models typically provide inherent additive feature attributions, assigning importance weights to individual or high-level features. For example,  Chen et al. (2019a) and  Wang & Wang (2021) describe model outputs by comparing them to relevant “prototypes”, while Alvarez Melis & Jaakkola (2018) derive concept coefficients from feature transformations. Other approaches, like those by Agarwal et al. (2021) and Jain et al. (2020), focus on feature-specific neural networks or apply classifiers to snippets for NLP explanations, respectively.  Koh et al. (2020) predict outcomes based on high-level concepts. Due to their focus on additive attributions, these methods generally operate under the implicit or explicit assumption that the model’s behavior can be approximated as linear under certain conditions (Yeh et al. (2019); Lundberg & Lee (2017)). This assumption is crucial as it underpins the approach of decomposing a model’s output into independent, additive contributions from individual features, and may hence overlook nonlinear interactions among features (Slack et al. (2020); Ribeiro et al. (2018); Yeh et al. (2019)). In contrast, we propose a self-explaining framework that provides a distinct form of explanation — sufficient reasons for predictions — which do not depend on assumptions of underlying linearity (Ribeiro et al. (2018)). To our knowledge, this is the first self-explaining framework designed to generate such explanations.
@@ 7 Limitations
A primary limitation of SST is the need to tune additional hyperparameters due to multiple optimization objectives. We simplified this by conducting a grid search solely over the cardinality loss coefficient $\xi$, as detailed in Appendix E. Additionally, the dual-propagation training phase is twice more computationally demanding than standard training, and this complexity escalates with challenging masking strategies, such as robust masking (see Appendix F for more details). However, this investment in training reduces the need for complex post-training computations, as highlighted in Section 5. Lastly, overly optimizing our interpretability constraints can sometimes reduce predictive accuracy. Our experiments showed this was true for IMAGENET, though all other benchmarks retained comparable accuracy levels (within $2\%$, and under $1\%$ for all language-model benchmarks). Interestingly, accuracy even improved under some configurations for MNIST classification. Future research endeavors could further investigate the conditions under which SST can be utilized not only for its inherent explanations but also to enhance predictive accuracy.
@@ Acknowledgments
  • Appendix A contains extended background on sufficient explanations and related work.
  • Appendix B contains background for the computational complexity proofs.
  • Appendix C contains the proof of theorem 1.
  • Appendix D contains the proof of theorem ii.
  • Appendix E contains technical specifications, related to the models, and training.
  • Appendix F contains information regarding the training time of SST compared to standard training.
  • Appendix G includes an experiment on the generalization of various masks across different sufficiency settings.
  • Appendix H includes an experiment on the extension of SST to simplified inputs.
  • Appendix I contains additional ablation results.
  • Appendix J contains supplementary results.
Appendix A Extended Related Work and Background
Sufficiency-based explanations. Sufficiency-based explanations are a widely sought-after approach in explainability, with numerous methods proposed to achieve them (e.g., (Ribeiro et al., 2018; Carter et al., 2019; Ignatiev et al., 2019; Bassan & Katz, 2023)). The core idea is to identify a minimal subset of features that determine a prediction, allowing one to focus on the essential “reason” for the outcome while excluding non-relevant features. Notably, unlike the more prevalent additive attribution methods (Ribeiro et al., 2016; Lundberg & Lee, 2017; Sundararajan et al., 2017), these explanations do not depend on approximate linearity assumptions (Ribeiro et al., 2018; Yeh et al., 2019). From a human perspective, prior research has demonstrated their effectiveness; for instance, seeing multiple sufficiency-based explanations helps users predict model decisions more accurately compared to relying on additive attributions (Ribeiro et al., 2018).
The concept of sufficient explanations has been extensively explored in the literature. The general notion of a sufficient explanation is closely linked to ideas from actual causality (Halpern, 2016) and has been examined in several works (Chockler et al., 2021; Chockler & Halpern, 2024; Chockler et al., 2024; Kelly et al., 2023; Watson et al., 2021). Furthermore, the robust class of sufficient explanations studied in this work also connects to research in formal logic (Marques-Silva, 2023; Darwiche & Marquis, 2021). In the broader case where the $\epsilon_{p}$ perturbation is unbounded, such explanations are sometimes referred to as abductive explanations (Ignatiev et al., 2019; Ordyniak et al., 2023; Bassan et al., 2023). A related but distinct concept is that of a prime implicant (Shih et al., 2018; Ignatiev et al., 2019). Given the high computational complexity of deriving such explanations (Barceló et al., 2020), significant research has focused on identifying simplified models where explanation computation remains tractable (Cooper & Marques-Silva, 2023; Audemard et al., 2021). Examples of such models include decision trees (Izza et al., 2020; Huang et al., 2021; Bounia & Koriche, 2023), KNNs (Barceló et al., 2025), tree ensembles (Izza & Marques-Silva, 2021; Ignatiev et al., 2022; Audemard et al., 2022b; a; Boumazouza et al., 2021), linear models (Marques-Silva et al., 2020; Subercaseaux et al., 2024), monotonic models (Marques-Silva et al., 2021; El Harzli et al., 2023), and small-scale neural networks (Ignatiev et al., 2019; Wu et al., 2024b; La Malfa et al., 2021; Bassan & Katz, 2023). These explanations are often facilitated by neural network verifiers (Wang et al., 2021b; Wu et al., 2024a; Fel et al., 2023).
Self-explaining models. The concept of a self-explaining neural network (SENN) was first introduced by (Alvarez Melis & Jaakkola, 2018) and advocates for training architectures that inherently provide interpretations for their decisions (Lee et al., 2022; Shwartz et al., 2020; Rajagopal et al., 2021; Guyomard et al., 2022; Guo et al., 2023; Zhang et al., 2022; Ji et al., 2025). This idea is closely related to other research on learning human-understandable “concepts” (Koh et al., 2020; Blazek & Lin, 2021; Espinosa Zarlenga et al., 2022; Marconato et al., 2022; Kim et al., 2023) and representative training-data “prototypes”(Chen et al., 2019a; Keswani et al., 2022; Hong et al., 2023). Additionally, related work explores training interventions designed to improve feature selection capabilities (Lemhadri et al., 2021; Zhang et al., 2024; Jethani et al., 2021; Jiang et al., 2024; Yang et al., 2022; Yamada et al., 2020).
Appendix B Computational Complexity Background